OpenStudy (anonymous):

A)Pick any five numbers. Compute the mean and standard deviation of this distribution. (List your five numbers.) Numbers: Mean: Standard Deviation: B) Multiply each number in your distribution by 3 and compute the mean and standard deviation of this new distribution. Numbers Now: Mean: Standard Deviation: C) Multiply each number in your original distribution by 9 and compute the mean and standard deviation of this new distribution. Numbers Now: Mean: Standard Deviation: D) What conclusion can you draw about the changes in the mean and the standard

myininaya (myininaya):

\[1,1,1,1,1\] mean=(1+1+1+1+1)/5

myininaya (myininaya):

\[variance=\frac{(1-1)^2+(1-1)^2+(1-1)^2+(1-1)^2+(1-1)^2}{5}=0\]

myininaya (myininaya):

\[standard deviation=\sqrt{variance}=\sqrt{0}=0\]

myininaya (myininaya):

i leave the rest to you

OpenStudy (anonymous):

ok..ty for the start

myininaya (myininaya):

i don't know if your teacher wanted such an easy data

OpenStudy (anonymous):

not sure its just confusing to me, im taking an online class

myininaya (myininaya):

if we had \[1,2,3,4,5\] \[mean=\frac{1+2+3+4+5}{5}=\frac{15}{3}=5\]

OpenStudy (anonymous):

I'm only doing part A but i'll show you how to do the rest. A) Numbers: 2, 4, 6, 8, 10 Mean(average): 2+4+6+8+10= 30 add together then divide by number of numbers (5) Mean=30/5 = 6 Standard Deviation = The square root of the average of the squared difference of the numbers from the mean. (mouthful?) \[SD=\sqrt{\sum(number-average)^{2}/number of numbers)}\] soo... (2-6)^2=16 (4-6)^2=4 (6-6)^2=0 (8-6)^2=4 (10-6)^2=16 Add those numbers together... = 40 standard deviation = sqrt{40/5}

myininaya (myininaya):

\[variance=\frac{(5-1)^2+(5-2)^2+(5-3)^2+(5-4)^2+(5-5)^2}{5}=\frac{16+9+4+1+0}{5}\]

myininaya (myininaya):

\[varinace=\frac{30}{5}=6\]

myininaya (myininaya):

\[standard deviation=\sqrt{variance}=\sqrt{6}\]

myininaya (myininaya):

there now you have three examples

myininaya (myininaya):

but i really like the data consisting of just 1's lol

OpenStudy (anonymous):

ty sooo much this is a big help!

OpenStudy (anonymous):

yea that's kind of funny

OpenStudy (anonymous):

online courses are nice in way but lac in examples and help